Parsons Plumbing & Heating manufactures thermostats that it uses in several of its products. Management is considering whether to continue manufacturing the thermostats or to buy them from an outside source. The following information is available.

1. The company needs 80,000 thermostats per year. Thermostats can be purchased from an outside supplier at a cost of $6 per unit.

2. The cost of manufacturing thermostats is $7.50 per unit, computed as follows.

Direct materials $ 156,000
Direct labor 132,000
Manufacturing overhead:
Variable 168,000
Fixed 144,000
Total manufacturing costs $ 600,000
Cost per unit ($600,000 ÷ 80,000 units) $ 7.50

  
3. Discontinuing the manufacture of the thermostats will eliminate all of the direct materials and direct labor costs but will eliminate only 60 percent of the variable overhead costs.

4. If the thermostats are purchased from an outside source, certain machinery used in the production process would no longer have to be leased. Accordingly, $9,200 of fixed overhead costs could be avoided. No other reductions will result from discontinuing production of the thermostats.

Required:

a-1. Prepare a schedule to determine the incremental cost or benefit of buying thermostats from the outside supplier.

a-2. On the basis of this schedule, would you recommend that the company manufacture thermostats or buy them from the outside source?

Assume that if thermostats are purchased from the outside source, the factory space previously used to produce thermostats can be used to manufacture an additional 6,000 heat-flow regulators per year. These regulators have an estimated contribution margin of $18 per unit. The manufacture of the additional heat-flow regulators would have no effect on fixed overhead.

b-1. Prepare a schedule showing the incremental cost or benefit of buying thermostats from the outside source and using the factory space to produce additional heat-flow regulators.

b-2. Would this new assumption change your recommendation as to whether to make or buy thermostats?
Answer #1

a1.

Manufacturing Purchasing
Direct Material 156000 0
Direct Labor 132000 0
Prime Cost 288000 0
Variable Manufacturing OH 168000 168000*(100-60)% = 67200
Fixed Manufacturing OH 144000 144000-9200 = 134800
Cost of Purchase 0 80000*6 = 480000
TOTAL 600000 682000

So, Incremental Cost => 682000-600000 = 82000$.

a2. yes, we recommend the company to manufacture than buy since it is cheaper than the latter.

.

b1.

Previously, costs under Purchase 682000
Contribution from regulators 6000*18 = -108000
Fixed overhead will be same as existing +9200
New Total Cost 583200

So, there would be Benefit of $ 16800 (ie, existing 600000 - 583200)

b2. Yes, we would recommend buying since it reduces the total costs.